The surface area s of a right prism is given by
s=2b+p h
,
where b is the area of the base, p is the perimeter of the base, and h is the height of the prism.
Solve for b .

Answer by RAY100(1637)

s = 2b + ph,,,,,,subtract ph both sides
.
s-ph = 2b,,,,,,,,,,divide both sides by 2
.
(s-ph)/2 =b
.
.
.
check,,,,s= 2(s-ph)/2 +ph = s-ph +ph =s,,,,,,,ok
.